# Statistics of correlations in nonlinear recurrent neural networks

### v1

@@ 5 Conclusions
In this work we have developed a path-integral framework to compute the statistics of correlations in nonlinear recurrent networks, including $1/N$ effects that are required to evaluate cross-correlations of the covariance matrix and the participaton dimension. Our approach generalizes previous treatments of linear networks by incorporating arbitrary nonlinear activation functions as interaction terms in the effective action. This resolves the instability of the linear theory, yielding strictly positive participation dimensions and a rich set of correlation structures.
A central outcome of our analysis is the emergence of a small set of collective variables that capture the full statistics of correlations. This representation makes explicit the way in which global network dynamics constrain local pairwise correlations. It also highlights the universality of some of the results: for instance, the $1/N$ suppression of cross-correlations persists in the nonlinear regime, but their relative fluctuations become essential in controlling the participation dimension. This is consistent with previous findings in the linear case [14, 15], while substantially extending their range of applicability. We also obtained the generating function of connected neural correlators, focusing for the most part on the covariance matrix; see Sec. 3.3. This is another key result of our work, and it is based on large $N$ factorization.
@@
Our results also allow us to relate the path-integral method to alternative theoretical frameworks. Dynamical mean-field theory and random matrix methods, which have been successfully applied to related questions [16, 17, 18], have provided valuable insights into the statistics of covariance matrices. The path-integral formulation complements these approaches by emphasizing the role of collective variables, the effects of external sources, and the description in terms of the generating function, which naturally encode higher-order correlators and enable systematic expansion in $1/N$. This complementary perspective could prove useful in situations where non-Gaussian features of the activity, or finite-size effects, play an essential role.
@@
In future work, it will be important to extend our approach to characterize out-of-equilibrium correlation functions, where temporal structure and nonstationary dynamics are expected to play a crucial role. Furthermore, the path-integral formulation provides an efficient framework to evaluate subleading $1/N$ effects using standard diagrammatic methods. For the recurrent networks studied here, such corrections originate from Gaussian fluctuations around the nontrivial saddle point as well as from interaction vertices induced by nonlinear activation functions. The path-integral framework is also promising for incorporating additional structures, such as plasticity. Exploring these contributions will deepen the connection between neural dynamics and field-theoretic techniques, and we expect to report on these developments in future work.
@@ Appendix A Correlations of neural inputs
$Z_{N_{r}}=\int d\phi\ d\eta d\rho\,\,e^{-\frac{i}{2}\eta_{ab}(\sum_{j}g_{ab}(\phi_{j})-N\rho_{ab})}\,[\textrm{det}(2\pi M^{-1})]^{N/2}\,e^{-\frac{1}{2}(G^{-1})_{ab}\phi_{i}^{a}\phi_{i}^{b}}e^{\sum_{j}J^{ab}_{j}\phi_{j}^{a}\phi_{j}^{b}}\,.$ (A.2)
@@
which should be compared with the analog expression (3.3) for the outputs. Here $U=U(G_{0})$ and $V=V(G_{0})$ are the functions defined in the main text, and $V^{\prime}=V^{\prime}(G_{0})$.
@@
Figure 5: Comparison of participation ratios of inputs and outputs, for $\beta=2,D=1$ (left) and $\beta=2,D=0.1$, right. Note the difference in scales in the y-axis between the plots.
The comparison between the analytical results and numerical simulations for the input correlations are shown in Figs. 6 and 7. As in the case of output correlations we fond an excellent agreement even for not very large network sizes.
Figure 6: Comparison of analytical and numerical results for the input correlation. Left panel: average value of the diagonal correlation. Central panel: average off-diagonal correlators. Right panel: participation ratio (participation dimension divided by network size). We use the transfer function of Eq.( 4.13) with $\beta=2$ and $D=1$. Network sizes $N=50,100,200$. The error bars represent the standard deviation over 5 simulations with different realizations of the coupling matrix $W$ and noise $\xi$.
Figure 7: Same as in the the previous figure but with the non saturating activation function of Eq. (4.15).
@@ Appendix B Network simulations
The differential equations that control the network dynamics, Eq.( 2.1), are simulated using a first order Euler method. As we intend to be in the limit $\tau\rightarrow 0$ at each time step we choose a vector of random Gaussian variables $\xi_{i}$ (1=1,N)(1=1,...N) with zero mean and variance $D$ and iterate
$\phi_{i}(t+\delta)=\phi_{i}(t)(1-\epsilon)+\left(\sum_{j=1}^{N}W_{ij}f(\phi_{i}(t))+\xi_{t}\right)\epsilon$ (B.1)
until the mean quadratic difference $d\equiv\frac{1}{N}\sum_{i=1}^{N}|\phi_{i}(t+\delta)-\phi_{i}(t)|^{2}$ is smaller than $10^{-4}$. We use $\delta=0.01$, $\epsilon=0.1$. This procedure is repeated $n_{l}=1000$ times to obtain the covariance matrices:
@@
where the brackets denote the temporal average: $x>>=\frac{1}{n_{l}}\sum_{l=1}^{n_{l}}x(l)$. As finite sampling generates bias in the statistics, specifically in the second order terms we implemented the corrections proposed in [19] (Supplementary Material).

### v2

@@ 4.2.2 Activation function with $p=1/2$
In the next section we present a brief discussion of related works and more specifically of the difference between our approch and the one in [32].
4.3 Comparison to other works and annealed vs quenched disorder
On the other hand, Ref. [11] focused on the linear network. In this work a path-integral formulation is also used. In contrast to our work there are no replicas and a linear source term is used. Correlations up to order four are evaluated taking succesive derivatives with respect to the source and the mean values and variances of the elements of the covariance matrix are determined using a Wick decomposition of the temporal averages. This is possible in the linear case because the statistics for fixed couplings is Gaussian. This approach is not possible in the non-linear case, requiring the use of replicas. Replicas have also been used in the linear model e.g. in [24].
Let us now analyze in more detail the difference between the quenched regime we have studied in this work, corresponding to $\tau/\tau_{\textrm{noise}}\ll 1$, and the annealed or white noise limit $\tau_{\textrm{noise}}/\tau\ll 1$, which has been studied in previous works, such as [19, 10]. In particular, the recent preprint [32] proposed an ansatz to solve the nonlinear network in terms of an effective coupling in the annealed limit. This was motivated by mean field theory considerations and verified numerically and using random matrix theory. Their ansatz can be derived explicitly from our large $N$ treatment. For this, we note that eq. (3.52) motivates the introduction of an effective coupling
$\lambda_{eff}\equiv\lambda U(G_{0})\,.$ (4.17)
$\lambda_{eff}=\lambda\int\frac{dx}{\sqrt{2\pi G}}\,f^{\prime}(x)\,e^{-\frac{x^{2}}{2G}}=\lambda\langle f^{\prime}(x)\rangle_{G}\,.$ (4.18)
This agrees with the effective connection strength Eq. (3.2) of [32].
As a further check of the consistency between both formalisms, we note that the output dimension of participation (3.52) becomes
$D_{PR}^{f}=N(1-\lambda^{2}\langle f^{\prime}(x)\rangle_{G}^{2})^{2}\,$, (4.19)
This coincides with Eq. (3.1) of [32] at vanishing frequency.
The main difference between the quenched and annealed limit lies in the time dependence of the neural correlators and their dependence on the model parameters. In the quenched limit and after reaching equilibrium, the neural correlators in equilibrium are time independent. Specifically, recalling the 2-point functions (3.31), we have that
$\langle C^{\phi}(t)\rangle_{W}=\langle\phi_{i}(t+t_{0})\phi_{i}(t_{0})\rangle_{\xi,\,W}\;,\;\langle C^{f}(t)\rangle_{W}=\langle f_{i}(\phi(t+t_{0}))f_{i}(\phi(t_{0}))\rangle_{\xi,\,W}\,$, (4.20)
are independent of $t$ and $t_{0}$, and coincide with what we have been calling $\langle C^{\phi}\rangle_{W}=G$ and $\langle C^{f}\rangle_{W}=GV(G)$ without time arguments. This implies that the frequency space correlators have only support at zero frequency:
$\langle C^{\phi}(\omega)\rangle_{W}=\int dt\,e^{-i\omega t}\,\langle C^{\phi}(t)\rangle_{W}=2\pi\delta(\omega)\,\langle C^{\phi}(t=0)\rangle_{W}$ (4.21)
and similarly for $C^{f}$. For completeness, let us also reproduce here the self-consistent equation (3.32),
$\langle C^{\phi}\rangle_{W}=D_{q}+\lambda^{2}\,\langle C^{f}\rangle_{W}\,$, (4.22)
where, for the following comparison with the annealed limit, we have renamed $D=D_{q}$ to highlight the quenched limit, and recall that
$\langle\xi_{i}(t+t_{0})\xi_{j}(t_{0})\rangle=\delta_{ij}D_{q}\,$, (4.23)
independent of time.
In the annealed limit, the self-consistent equation for the frequency space neural correlators is, from Eq. (2.2) of [32],
$(1+\omega^{2})\langle C^{\phi}(\omega)\rangle_{W}=D_{w}+\lambda^{2}\,\langle C^{f}(\omega)\rangle_{W}\,$, (4.24)
where the white noise coupling $D_{w}=\sigma^{2}$ (in their notation) satisfies
$\langle\xi_{i}(t+t_{0})\xi_{j}(t_{0})\rangle=\delta_{ij}\delta(t)D_{w}\,.$ (4.25)
The quenched and annealed noise parameters $D_{q}$ and $D_{w}$ cannot be compared directly, as they have different dimensions. However, we can obtain a more precise relation between (4.22) and (4.24) if instead of $D_{w}$ we write
$D_{w}=\int dt\,e^{-i\omega t}\,\langle\xi_{i}(t)\xi_{i}(0)\rangle=\langle C^{\xi}(\omega)\rangle\,$, (4.26)
and propose a self-consistent equation
$(1+\omega^{2})\langle C^{\phi}(\omega)\rangle_{W}=\langle C^{\xi}(\omega)\rangle+\lambda^{2}\,\langle C^{f}(\omega)\rangle_{W}\,.$ (4.27)
We propose that this self-consistent equation is valid for a more general internal “colored noise”, corresponding to $\tau_{\textrm{noise}}$ not parametrically different from $\tau$. We hope to check the consequences of this in future work. It correctly interpolates between the white and quenched limits. In the white noise limit, (4.27) reduces to (4.24), while in the quenched limit it reduces to (4.22) since all correlators are proportional to $\delta(\omega)$.
To have a better understanding of the theoretical limits and the parameter ranges, let us now compare numerical results in the quenched and white noise limits. In terms of the neural time constant $\tau$, it is natural to contrast results for $D_{w}\sim D_{q}/\tau$, which have the same time dimensions. As a concrete example, let us choose
$D_{w}=\frac{D_{q}}{2\tau},$ (4.28)
which corresponds to a situation where the equal time neural 2-point function in a linear system with white noise agrees with the quenched equilibrium 2-point function. [^8] Again, this is just an example and there is no fundamental reason for such a relation between $D_{w}$ and $D_{q}$, especially in nonlinear cases. In Fig 6 we show the numerical results for $N=100,\tau=1,D_{w}=2$ and $D_{q}=1$ with other parameters as in Fig. 5. This shows a qualitative similarity between both results in the nonlinear case. We see that even though the quenched and white noise models are quite different mathematical idealizations, they provide complementary results which may not be that different from more experimentally realistic models.
Figure 6: Comparison of annealed and quenched dynamics with $\tau=1$ for the output correlation for the Padé activation function with $p=1/2$, $N=100,\beta=2$ and $D=2(1)$ for the annealed (quenched) case.
@@ 5 Conclusions
In this work we have developed a path-integral framework to compute the statistics of correlations in nonlinear recurrent networks, including $1/N$ effects that are required to evaluate cross-correlations of the covariance matrix and the participaton dimension. Our approach generalizes previous treatments of linear networks by incorporating nonlinear activation functions as interaction terms in the effective action. [^9] For simplicity we restricted to odd activation functions, which lead to a vanishing neural 1-point function. More general activation functions can be included by allowing for such an expectation value. This resolves the instability of the linear theory, yielding strictly positive participation dimensions and a rich set of correlation structures.
A central outcome of our analysis is the emergence of a small set of collective variables that capture the full statistics of correlations. This is a basic result in large $N$ and disordered average field theories, as well as in the case of linear networks; here we have extended the collective representation to nonlinear networks. This representation makes explicit the way in which global network dynamics constrain local pairwise correlations. It also highlights the universality of some of the results: for instance, the $1/N$ suppression of cross-correlations persists in the nonlinear regime, but their relative fluctuations become essential in controlling the participation dimension. This is consistent with previous findings in the linear case [11, 19], while substantially extending their range of applicability. We also obtained the generating function of connected neural correlators, focusing for the most part on the covariance matrix; see Sec. 3.3.
@@
Our results have been derived in the quenched disordered limit, and we have compared to alternative frameworks that use white noise (annealed internal disorder). Dynamical mean-field theory and random matrix methods, which have been successfully applied to related questions [7, 36, 32], have provided valuable insights into the statistics of covariance matrices. The path-integral formulation complements these approaches by emphasizing the role of collective variables, the effects of external sources, and the description in terms of the generating function, which naturally encode higher-order correlators and enable systematic expansion in $1/N$. Both the quenched and annealed cases provide mathematically tractable regimes, and can lead to comparable predictions. This suggests that more realistic colored noise models may also behave similarly. In fact, based on the comparison between both limits, we have proposed a new self-consistent equation (4.27) for the generic colored noise case. In future work we plan to derive this from a large $N$ treatment, and study its dynamical consequences.
@@
In future work, it will be important to extend our approach to characterize out-of-equilibrium correlation functions, where temporal structure and nonstationary dynamics are expected to play a crucial role. Furthermore, the path-integral formulation provides an efficient framework to evaluate subleading $1/N$ effects using standard diagrammatic methods. For the recurrent networks studied here, such corrections originate from Gaussian fluctuations around the nontrivial saddle point as well as from interaction vertices induced by nonlinear activation functions.
It is interesting that in the quenched regime that our work has addressed, the large $N$ limit seems to work extremely well already for moderate values $N\sim 100$, while the annealed case requires much larger values of NN. [^10] We thank an annonymous referee for stressing this. We have performed a numerical study of the $1/N$ convergence, but it would be interesting to have a detailed analytic understanding of the large $N$ regime of validity. This will require going beyond 4-point functions as well as including “fluctuation corrections” from perturbations around the saddle points. Finally, the path-integral framework is also promising for incorporating additional structures, such as plasticity. Exploring these contributions will deepen the connection between neural dynamics and field-theoretic techniques, and we expect to report on these developments in future work.
@@ Appendix A Correlations of neural inputs
$Z_{N_{r}}$ $=\int{\mathcal{D}}\phi\,{\mathcal{D}}\eta\,{\mathcal{D}}\rho\;\exp\!\left[-\frac{i}{2}\sum_{a,b=1}^{N_{r}}\eta_{ab}\Big(\sum_{j=1}^{N}g_{ab}(\phi_{j})-N\rho_{ab}\Big)\right]$
$\times[\textrm{det}(2\pi M^{-1})]^{N/2}\,\exp\!\left[-\frac{1}{2}\sum_{j=1}^{N}\sum_{a,b=1}^{N_{r}}(G^{-1})_{ab}\phi_{j}^{a}\phi_{j}^{b}\right]\,\exp\!\left[\sum_{j=1}^{N}\sum_{a,b=1}^{N_{r}}J^{ab}_{j}\phi_{j}^{a}\phi_{j}^{b}\right].$ (A.2)
@@
which should be compared with the analog expression (3.3) for the outputs.
Here $U=U(G_{0})$ and $V=V(G_{0})$ are the functions defined in the main text, and $V^{\prime}=V^{\prime}(G_{0})$.
@@
The comparison between the analytical results and numerical simulations for the input correlations are shown in Figs. 8. As in the case of output correlations we found an excellent agreement even for not very large network sizes.
Figure 7: Comparison of participation ratios of inputs and outputs, for $\beta=2,D=1$ (left) and $\beta=2,D=0.1$ (right), with a Pade activation function $p=0$ (4.13). Note the difference in scales in the y-axis between the plots.
Figure 8: Comparison of analytical and numerical results for the input correlation. Left column: average values of the diagonal correlations. Central column: average off-diagonal correlators. Right column: participation ratio (participation dimension divided by network size). In the upper row we use the transfer function of Eq.(4.1) with $\alpha=1$, $p=1/2$ and $D=1$; in the center we use the transfer function of Eq.( 4.13) with $\beta=2$ and $D=1$, and in the lower row we use the non saturating activation function of Eq. (4.15). Network sizes $N=50,100,200$. The error bars represent the standard deviation over 5 simulations with different realizations of the coupling matrix $W$ and noise $\xi$.
@@ Appendix B Network simulations
The differential equations that control the network dynamics in the quenched limit, Eq. (2.7), are simulated using a first order Euler method. As we intend to be in the limit $\tau\rightarrow 0$ at each time step we choose a vector of random Gaussian variables $\xi_{i}$ (1=1,N)(1=1,...N) with zero mean and variance $D$ and iterate
$\phi_{i}(t+\delta)=\phi_{i}(t)(1-\epsilon)+\left(\sum_{j=1}^{N}W_{ij}f(\phi_{j}(t))+\xi_{i}\right)\epsilon$ (B.1)
until the mean quadratic difference $d\equiv\frac{1}{N}\sum_{i=1}^{N}|\phi_{i}(t+\delta)-\phi_{i}(t)|^{2}$ is smaller than $10^{-4}$. We use $\delta=0.01$, $\epsilon=0.1$. This procedure is repeated $n_{t}=1000$ times to obtain the covariance matrices:
@@
where the brackets denote the trial average: $x>>=\frac{1}{n_{t}}\sum_{l=1}^{n_{t}}x(l)$. As finite sampling generates bias in the statistics, specifically in the second order terms, we implemented the corrections proposed in [11] (Supplementary Material).
The results shown in Section 4.3 for the opposite limit ($\tau_{noise}\ll\tau$) were obtained by simulating Eq.(2.1) using the Euler-Maruyama method, i.e.
$\phi_{i}(t+\delta)=\phi_{i}(t)\left(1-\frac{\delta}{\tau}\right)+\left(\sum_{j=1}^{N}W_{ij}f(\phi_{j}(t))\right)\frac{\delta}{\tau}+\frac{\sqrt{\delta}}{{\tau}}\xi_{i}$ (B.4)
where $\xi_{i}$ is a vector of random variables with Gaussian distribution (zero mean and variance $D$) that are extracted independently at each time step. The results of the simulations were used to evaluate the time lagged correlation functions in the same way as in [32].
In Fig. 9 we show representative traces for the two dynamics. In the upper panels we display the convergence of the quenched dynamics (Eq. (2.7)) to the fixed point determined by the values of the noise variable. In the lower panel we show the very fast changes induced by the annealed dynamics of Eq. (2.1).
Figure 9: Comparison of dynamical behavior for the quenched (upper panels) and annealed dynamics (lower panels). Sample traces of 5 neurons for different values of $\lambda$. Paremeters as in Fig. 2 with $N=100$.
Appendix C Derivation of the effective action at large $N$ at order $J^{2}$
In this Appendix for completeness we present the details of the derivation of the effective action (3.7).
$\hskip-85.35826pt=\langle e^{-\tfrac{i}{2}{\rm tr}(\eta g)}\rangle_{G}\,\textrm{exp}\!\left[\frac{\langle e^{-\tfrac{i}{2}{\rm tr}(\eta g)}\,{\rm tr}\!\left(J_{j}(g+xx)\right)\rangle_{G}}{\langle e^{-\tfrac{i}{2}{\rm tr}(\eta g)}\rangle_{G}}\right]$ (C.1)
$\langle\prod_{j}e^{-\tfrac{i}{2}{\rm tr}(\eta g)+{\rm tr}\left(J_{j}(g+xx)\right)}\rangle_{G}$ $=\langle e^{-\tfrac{i}{2}{\rm tr}(\eta g)}\rangle^{N}_{G}\,\textrm{exp}\!\left[\sum_{j}\frac{\langle e^{-\tfrac{i}{2}{\rm tr}(\eta g)}\,{\rm tr}\!\left(J_{j}(g+xx)\right)\rangle_{G}}{\langle e^{-\tfrac{i}{2}{\rm tr}(\eta g)}\rangle_{G}}\right]$ (C.2)
As discussed in the main text, we focus for concreteness on calculating the partition function including second order contributions from the sources. Terms at higher orders an be obtained following the same procedure. We write this as
$Z_{N_{r}}(J)=\int{\mathcal{D}}\eta\,{\mathcal{D}}\rho\,e^{-NS_{eff}}\,.$ (C.3)
Inserting (C.2) into (3.5) and collecting all factors in the exponential, we obtain, up to $\mathcal{O}(J^{2})$,
$S_{\mathrm{eff}}$ $=\frac{1}{2}\log\det(1+iM\eta)-\frac{i}{2}{\rm tr}(\eta\rho)-\log\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g(x))}\Big\rangle_{G}$
$\quad-\frac{1}{N}\sum_{j}\frac{\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g(x))}\,{\rm tr}\!\big(J_{j}(g(x)+xx)\big)\Big\rangle_{G}}{\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g(x))}\Big\rangle_{G}}$
$\quad-\frac{1}{2N}\sum_{j}\frac{\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g(x))}\,\big({\rm tr}\!\big[J_{j}(g(x)+xx)\big]\big)^{2}\Big\rangle_{G}}{\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g(x))}\Big\rangle_{G}}$
$\quad+\frac{1}{2N}\sum_{j}\frac{\Big(\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g(x))}\,{\rm tr}\!\big(J_{j}(g(x)+xx)\big)\Big\rangle_{G}\Big)^{2}}{\Big(\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g(x))}\Big\rangle_{G}\Big)^{2}}+\ldots$ (C.4)
The cumulant expansion of the first expectation value gives
$-\log\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g)}\Big\rangle_{G}$ $=\frac{i}{2}{\rm tr}\!\big(\eta\langle g\rangle_{G}\big)+\frac{1}{8}\Big(\big\langle({\rm tr}(\eta g))^{2}\big\rangle_{G}-\big\langle{\rm tr}(\eta g)\big\rangle_{G}^{\,2}\Big)+\mathcal{O}(\eta^{3}).$ (C.5)
Note that (C.2) is a cumulant expansion in the sources $J_{j}$ (keeping the $\eta$-dependent weight exact), whereas (C.5) follows from a separate cumulant expansion in $\eta$. Using the symmetry of $\eta$ and $g$,
${\rm tr}(\eta g)=\sum_{a}\eta_{aa}g_{aa}+2\sum_{a<b}\eta_{ab}g_{ab},$ (C.6)
and expanding the $\eta$-dependence:
$-\log\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g)}\Big\rangle_{G}$ $\approx\frac{i}{2}{\rm tr}\!\big(\eta\langle g\rangle_{G}\big)+\frac{1}{2}\sum_{a<b}\eta_{ab}^{2}\,\langle g_{ab}^{2}\rangle_{G}.$ (C.7)
For the linear source contribution we expand the ratio in (C.4) to first order in $\eta$:
$\frac{\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g)}\,{\rm tr}\!\big(J_{j}(g+xx)\big)\Big\rangle_{G}}{\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g)}\Big\rangle_{G}}$ $=\Big\langle{\rm tr}\!\big(J_{j}(g+xx)\big)\Big\rangle_{G}-\frac{i}{2}\Big\langle{\rm tr}(\eta g)\,{\rm tr}\!\big(J_{j}(g+xx)\big)\Big\rangle_{G}$
$\quad+\frac{i}{2}\Big\langle{\rm tr}(\eta g)\Big\rangle_{G}\Big\langle{\rm tr}\!\big(J_{j}(g+xx)\big)\Big\rangle_{G}+\mathcal{O}(\eta^{2}J),$ (C.8)
To obtain an explicit expression in replica indices, we write ${\rm tr}\!\big(J_{j}(g+xx)\big)=\sum_{ab}J_{j}^{ab}\big(g_{ab}+x^{a}x^{b}\big)$ and use $\langle x^{a}x^{b}\rangle_{G}=G_{ab}$. We then separate diagonal and independent off-diagonal components in replica space and keep only the contributions that survive at the leading orders under the large-$N$ scaling stated above. In particular, in the $\mathcal{O}(\eta J)$ term we retain the off-diagonal sector $a\neq b$ and the leading pairwise contractions at large $N$, which yields
$\frac{\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g)}\,{\rm tr}\!\big(J_{j}(g+xx)\big)\Big\rangle_{G}}{\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g)}\Big\rangle_{G}}$ $=\sum_{a}J_{j}^{aa}\Big(G_{aa}+\langle g_{aa}\rangle_{G}\Big)+\sum_{a<b}J_{j}^{ab}\Big(G_{ab}+\langle g_{ab}\rangle_{G}\Big)$
$\quad-i\sum_{a<b}\eta_{ab}\,J_{j}^{ab}\,\Big\langle g_{ab}\big(g_{ab}+x^{a}x^{b}\big)\Big\rangle_{G}+O(\eta^{2}J).$ (C.9)
For the quadratic source contribution we only need the leading term at $\eta=0$, since any $\eta$-dependence would produce subleading corrections of order $O(\eta J^{2})$ under the same scaling:
$\frac{\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g)}\,\big({\rm tr}\!\big[J_{j}(g+xx)\big]\big)^{2}\Big\rangle_{G}}{\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g)}\Big\rangle_{G}}-\frac{\Big(\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g)}\,{\rm tr}\!\big(J_{j}(g+xx)\big)\Big\rangle_{G}\Big)^{2}}{\Big(\Big\langle e^{-\frac{i}{2}{\rm tr}(\eta g)}\Big\rangle_{G}\Big)^{2}}$
$\qquad=\sum_{a}(J_{j}^{aa})^{2}\Big(\big\langle(g_{aa}+x^{a}x^{a})^{2}\big\rangle_{G}-\big\langle g_{aa}+x^{a}x^{a}\big\rangle_{G}^{2}\Big)+\sum_{a<b}(J_{j}^{ab})^{2}\,\big\langle(g_{ab}+x^{a}x^{b})^{2}\big\rangle_{G}+O(\eta J^{2}),$ (C.10)
We have checked self-consistently that the diagonal components of $\rho_{ab}$ and $G_{ab}$ are order $N^{0}$, while the off-diagonal components of $\rho_{ab}$ and $G_{ab}$, as well as all the variables $\eta_{ab}$ are order $1/N$. With this $N$-scaling, the leading terms at large $N$ are
$S_{eff}$ $\approx\frac{1}{2}\log\det(1+iM\eta)-\frac{i}{2}{\rm tr}(\eta\rho)+\frac{i}{2}{\rm tr}\left(\eta\langle g\rangle_{G}\right)+\frac{1}{2}\sum_{a<b}\eta_{ab}^{2}\langle g_{ab}^{2}\rangle_{G}$ (C.11)
$-\frac{1}{2N}\sum_{j}\left[\sum_{a}(J_{j}^{aa})^{2}\left(\langle(g_{aa}+x^{a}x^{a})^{2}\rangle_{G}-\langle g_{aa}+x^{a}x^{a}\rangle_{G}^{2}\right)+\sum_{a<b}(J_{j}^{ab})^{2}\langle(g_{ab}+x^{a}x^{b})^{2}\rangle_{G}\right]\,$,
which reproduces (3.7).